how do i solve this problem?
Gigi's kids love honey-sweetened cerel. She mixes the hone-sweetened cereal with the unsweetened version of the same cereal. From experience, she knows that the misture is perfect whenthe amount of honey-sweetened cereal is exactly 40% of the mix. Her husband, Chase, has just mixed together a 14oz package of unsweetened cereal and a 32oz package of sweetened cereal. how much more unsweetened cereal does Gigi need to mix in to make th emix 40% sweetened and 60% unsweetened?
(i would show what i worked on, but i don't know how to start it and what to do.)

Answer by ankor@dixie-net.com(22740)

From experience, she knows that the mixture is perfect when the amount of honey-sweetened cereal is exactly 40% of the mix. Her husband, Chase, has just mixed together a 14 oz package of unsweetened cereal and a 32 oz package of sweetened cereal. how much more unsweetened cereal does Gigi need to mix in to make the mix 40% sweetened and 60% unsweetened?
:
Dear student,
I made this problem unnecessarily complicated, think of it this way:
How much much unsweetened cereal will we have add to the original mixture to
get 40%?
:
Original mixture amt: 14 + 32 = 46 oz
:
We have to add the amt that will make 32 oz equal to 40% of the final mixture.
:
Let x = amt of unsweetened cereal to be added
:
32 = .40(46 + x)
:
32 = 18.4 + .4x
:
32 - 18.4 = .4x
:
13.6 = .4x
:
x = 13.6/.4
:
x = 34 oz of unsweetened cereal required
:
Check our solution:
Total amt after adding the unsweetened stuff: 34 + 46 = 80 oz
sweet cereal percentage: 32/80 = .4; which is 40% as required
:
Did this make MORE sense to you now?